# Multi-perspective Alignment for Increasing Naturalness in Neural Machine Translation

### v1

@@ Abstract
Huiyuan Lai1, Esther Ploeger2, Rik van Noord1, Antonio Toral1 1CLCG, University of Groningen, The Netherlands 2Department of Computer Science, Aalborg University, Denmark {h.lai,r.i.k.van.noord,a.toral.ruiz}@rug.nl     espl@cs.aau.dk
@@ 1 Introduction
While machine translation (MT) has achieved promising performance with the adoption of neural networks (Bahdanau et al., 2015; Vaswani et al., 2017; Team NLLB et al., 2022), automatic translations remain markedly different from translations by professional human translators. A striking example is the fact that MT outputs exhibit reduced lexical diversity (Vanmassenhove et al., 2019, 2021) and increased source-language interference (Toral, 2019) compared to human translation (HT). These linguistic differences were previously referred to as machine translationese (de Clercq et al., 2020; Bizzoni et al., 2020; Vanmassenhove et al., 2021). [^1] This term has since been criticized, see for example Crespo (2023).
@@
To address these challenges, we frame the task of increasing naturalness in MT as a text style transfer-like task, where style and content are the two core aspects (Mou and Vechtomova, 2020; Lai et al., 2021). We train a vanilla MT model with supervised learning and subsequently exploit reward learning that fosters naturalness and content preservation. With respect to naturalness we explore two objectives: making MT more akin to human translations (i.e. reducing machine translationese) and making MT more akin to texts originally written in the target language, i.e. reducing translationese (Gellerstam, 1986; Baker, 1993; Toury, 2012). We evaluate our framework on a dataset for English-to-Dutch literary translation. Our main contributions are as follows:
@@
  • Extensive experiments show that our model produces translations that are lexically richer than baseline MT systems without loss in translation quality.
@@ 2.1 Increasing MT Naturalness
A few approaches have been put forward to make NMT output more natural. For example, Freitag et al. (2019) trained a post-processor that learns to translate from round-trip machine translated to original text in the same language. Freitag et al. (2022) prepend their training examples with special tags that denote whether the target side of the training data was originally written in that language or not. These methods are rigid, while in some cases, content preservation may be more important than style (Parthasarathi et al., 2021). In response, Ploeger et al. (2024) propose a flexible approach based on reranking translation candidates, but report considerable loss in general translation quality. Our method is tailorable to the downstream scenario, while still being faithful to the source texts.
A slightly related line of work aims to reduce translationese from human translations, and uses monolingual approaches based on style transfer (Jalota et al., 2023), semantic parsing (Wein and Schneider, 2024) and debiasing embeddings (Dutta Chowdhury et al., 2022). The other related line is to leverage human feedback to improve overall translation quality where a single metric such as COMET trained from human annotations is used as the reward model (Ramos et al., 2024; He et al., 2024). In this work we aim to improve translation quality from multiple perspectives.
@@
Baroni and Bernardini (2005) showed that original texts can be distinguished from human-translated texts with computational methods. Concrete textual markers, such as the frequency of function words or the use of punctuation, have been associated with this difference (Koppel and Ordan, 2011; Volansky et al., 2015). Beyond hand-crafting specific linguistic features, Pylypenko et al. (2021) find that neural architectures provide a reliable tool for distinguishing translated from original texts. They obtain best performance by fine-tuning multilingual BERT (Devlin et al., 2019) on the task, retrieving average accuracies ranging from 84.6% to 94.4%.
@@ 3 Data
We use the parallel dataset from Toral et al. (2021), preprocessed by Ploeger et al. (2024). This dataset consists of 531 books that were originally written in English (OR) and human translated into Dutch (HT), of which 495 books for training, 5 for validation and 31 as a test set. The genres of these books vary, including literary fiction, popular fiction, non-fiction and children’s books from over 100 authors. Particularly, the test set also contains a broad range of books. [^2] A full list of author names, titles, genres and publishing years of the test set books can be found in the Appendix. In addition, we use monolingual data for the two baseline MT systems (see Section 5.1), consisting of a random sample of equal size to the parallel training data and disjoint from the subset used for translation detection.
@@ 4.1 Base MT Model
For HT vs OR classification, we use the monolingual Dutch data introduced in the first part of Table 1. For the other two settings, we translated a subset of the English text in the parallel data (second part of Table 1) of equal size to the monolingual training data (982,114 sentences) into Dutch using the base MT model. The resulting machine translated sentences are combined with OR texts in the monolingual data for MT vs OR classification and with HT texts in the parallel data for MT vs HT. We filter out machine translated texts that are identical to human translations.
Based on the above data, we fine-tune the Dutch language model BERTje (de Vries et al., 2019) for the binary detection tasks, obtaining three different models. On the test sets, they achieve an accuracy of 84.8% on HT vs OR, 78.2% on MT vs HT, and 91.1% on MT vs OR. This is on par with the performance in similar scenarios from previous work (Pylypenko et al., 2021). Also in line with previous work (van der Werff et al., 2022), we find that the distinction between the translation variants (MT and HT) is especially challenging.
Algorithm 1 Multi-perspective alignment algorithm for translationese reduction
@@
1:  for each iteration $i=0,1,\cdots,m$ do
@@ 4.3 Multi-perspective Alignment for Naturalness and Content Preservation
We introduce our method which ranks samples based on rewards that target naturalness and content preservation. This approach is inspired by recent work in text style transfer, where both meaning has to be preserved and style should be transferred (Lai et al., 2021). This content vs form trade-off is similar to our situation with content preservation and naturalness. Specifically, after training a base MT model using supervised learning (Section 4.1), we further align it with human expectations in terms of naturalness and content in the form of reward learning.
@@
where $\phi$ is the parameter of the classifier. $\sigma_{t}$ is the translationese threshold (set to 0.5 in our experiments).
@@
Where C($\cdot$) represents the COMET model and $\sigma_{t}$ represents the content threshold (set to 0.85 in our experiments).
@@
To keep the fine-tuned model from moving too far from the base MT model, we combine the reward objective with the supervised training loss instead of using a reference model requiring large computing resources. Therefore, the final objective function of our framework consists of two components: negative log-likelihood loss in Eq. 1 and reward-based loss in Eq. 5, jointly formulated as
@@ 5.1 Baselines
In addition to the base MT model (Section 4.1), we include two previous methods that aim at reducing machine translationese as baselines: automatic post-editing (APE) (Freitag et al., 2019) and Tagging (Freitag et al., 2022).
@@ 5.2 Implementation Details
All experiments are implemented using the library HuggingFace Transformers (Wolf et al., 2020). We use the BART (Lewis et al., 2020) architecture with 6 Transformer-based (Vaswani et al., 2017) layers in both the encoder and decoder. The base MT models are trained using the AdamW optimiser (Loshchilov and Hutter, 2019) with a cosine learning rate decay, and a linear warmup of 1,000 steps. The maximum learning rate is set to 1e-4, the batch size is 256, and the gradient accumulation is 2; all reward-based models are trained with a consistent learning rate of 2e-5. We evaluate the model every 1,000 steps and use early stopping with patience 3 if the cross-entropy loss on the validation set does not decrease. We use beam search with size 5 during inference. Since some of the training data contains instances of repeated punctuation marks, this led to the reinforcement learning method tending to optimize the model for higher rewards. Therefore, we take a simple post-processing step to remove consecutive repeated punctuation marks after the text is generated. [^3] See Appendix A.2 for post-processing examples.
@@ 5.3 Evaluation Methods
We use three metrics to automatically calculate the content preservation of the output based on human references (and source sentences): BLEU (Papineni et al., 2002), COMET (Rei et al., 2020, 2022), and MetricX (Juraska et al., 2024). We use the Sacre-BLEU implementation (Post, 2018) for BLEU. Regarding the COMET family, we use both the default model wmt22-comet-da (COMET), and the reference-free model wmt22-cometkiwi-da (KIWI) that is not used for reward learning. For MetricX, we use MetricX-24-Hybrid-XL, considering it our most important translation quality metric, since it achieved state-of-the-art performance on the WMT24 Metrics Shared Task (Freitag et al., 2024).
Figure 2: Evaluation results under various settings. Notes: (i) The iteration step of 0K represents the base MT model; HM indicates the harmonic mean of classification accuracy (i.e. HT-OR, MT-HT or MT-OR) and COMET score.
@@
We employ the translationese detection models to assess outputs and report the classification accuracy. Additionally, as previous studies show that translated texts are often simpler than original texts (Baker, 1993), our evaluation also covers lexical diversity. Here we report six different metrics:
@@
  • MTLD (McCarthy, 2005): evaluated sequentially as the average length of sequential word strings in a text that maintains a given TTR value. We use a threshold of 0.72.
  • B1 (Vanmassenhove et al., 2021): the percentage of words that belong to the 1,000 most frequent words.
@@ 6.1 Initial Results
During the alignment training phase (see Section 4.3) we find that the loss does not correlate with MT quality, especially in terms of the naturalness aspect (i.e. classifier’s accuracy): while naturalness improves, the loss on the validation set stays flat. Therefore we manually select checkpoints between the 1k and the 6k training steps, and report their evaluation curves in Figure 2.
The first observation is that all models achieve substantial improvement in naturalness over the first 1k steps compared to the base MT model (i.e. 0K), as reflected in the results for translationese classification (HT-OR, MT-HT and MT-OR) and lexical richness (MTLD). Although the COMET scores of some models decrease slightly, the overall score HM follows the trend of the translationese aspect. After 1k steps, MTLD scores tend to be flat, and the translationese classification has some fluctuations but overall keeps improving throughout. For the remaining experiments, we report the results of the alignment model at 5K iteration steps.
@@
Table 3: Ablation study: Evaluate the contribution of each reward component, where we fine-tune the base MT model using only the content reward or the naturalness reward.
6.2 Main Results
We report the main evaluation results in Table 2, including the base MT model, the two baselines and our methods trained with both rewards: COMET for content preservation and the three different classifiers for naturalness (i.e. HT-OR, MT-HT and MT-OR).
Compared to APE, Tagging consistently performs better across the board, both in terms of content (i.e. translation accuracy) and naturalness. Additionally, we observe that using more target-original data results in lower accuracy scores but better naturalness metrics, which is consistent with our hypothesis (see Section 5.1). We also observe that both baselines underperform the base MT model in terms of translation accuracy.
When comparing different classification rewards, the model trained with COMET & MT-HT achieves, overall, better scores than our other two models (HT-OR and MT-OR). We speculate that the rewards that foster OR do not work as well due to a mismatch between the preference of the classifier (OR) and the data in the target side of the MT training data (HT). We thus speculate that such classifiers could be useful in scenarios in which the target side of the MT training data contains texts originally written in that language, which would be common in translation directions in which the target language is higher-resourced than the source language.
Overall, our best system (BM + COMET & MT-HT) achieves better naturalness scores than the base MT model (e.g. 93.3 vs 90.4 for MTLD), while even having a higher KIWI score (80.6 vs 80.4) and a lower MetricX score (2.63 vs 2.66; lower scores are better), two metrics that have not yet been used for reward learning. Tagging attains higher naturalness scores but this comes at the price of a notable reduction in translation accuracy, as shown by KIWI (79.7 vs 80.6) and MetricX (3.05 vs 2.63).
@@ 6.3 Ablation Study
To assess the contribution of each reward component, we perform a set of ablation studies, as shown in Table 3. For the COMET vs COMET + classifier setting, we see higher naturalness scores in the latter in all cases for MT-HT and MT-OR (except CDU in MT-HT), as expected, while there are mixed cases in HT-OR. Also as expected, translation accuracy scores decrease when the naturalness reward is added (except COMET with MT-HT).
Compared to classifier-only models, classifier + COMET results generally in better naturalness-related metrics (except PTF), but worse content-based metrics (except COMET with MT-HT). This might be due to a mismatch between the classifier’s objective and the training data (see comment in Section 6.2) and to complex interactions between both rewards, that would require further inspection.
@@
Table 4: Example of human-written text (source and HT), the most relevant baselines (Tagging, BM) and ours.

### v2

@@ Abstract
Huiyuan Lai1, Esther Ploeger2, Rik van Noord1, Antonio Toral3thanks: Work partly carried out while affiliated with the University of Groningen. 1CLCG, University of Groningen, The Netherlands 2Department of Computer Science, Aalborg University, Denmark 3Departament de Llenguatges i Sistemes Informàtics, Universitat d’Alacant, Spain {h.lai,r.i.k.van.noord}@rug.nl     espl@cs.aau.dk     antonio.toralr@ua.es
@@ 1 Introduction
While machine translation (MT) has achieved promising performance with the adoption of neural network (Bahdanau et al., 2015; Vaswani et al., 2017; Team NLLB et al., 2022), automatic translations remain markedly different from translations by professional human translators. A striking example is the fact that MT outputs exhibit reduced lexical diversity (Vanmassenhove et al., 2019, 2021) and increased source-language interference (Toral, 2019) compared to human translation (HT). These linguistic differences were previously referred to as machine translationese (de Clercq et al., 2020; Bizzoni et al., 2020; Vanmassenhove et al., 2021). [^1] This term has since been criticized, see for example Crespo (2023).
@@
To address these challenges, we frame the task of increasing naturalness in MT as a text style transfer-like task, where style and content are the two core aspects (Mou and Vechtomova, 2020; Lai et al., 2021b, a). In practice, we train a vanilla MT model with supervised learning and subsequently exploit reward learning that fosters naturalness and content preservation, as shown in Figure 1. With respect to naturalness we explore two objectives: making MT more akin to human translations (i.e. reducing machine translationese) and making MT more akin to texts originally written in the target language, i.e. reducing translationese (Gellerstam, 1986; Baker, 1993; Toury, 2012). We evaluate our framework on a dataset for English-to-Dutch literary translation. Our main contributions are as follows:
@@
  • Extensive experiments show that our model produces translations that are lexically richer than baseline MT systems without loss in translation quality. [^2] All code at https://github.com/laihuiyuan/alignment4naturalness
@@ 2.1 Increasing MT Naturalness
A few approaches have been put forward to make MT outputs more natural. For example, Freitag et al. (2019) trained a post-processor that learns to translate from round-trip machine translated text to original text in the same language, which can be applied to the outputs of existing MT systems. Freitag et al. (2022) prepend their training examples with special tags that denote whether the target side of the training data was originally written in that language or not. These methods are rigid, while in some cases, content preservation may be more important than style (Parthasarathi et al., 2021). In response, Ploeger et al. (2024) propose a flexible approach based on reranking translation candidates, but report considerable loss in general translation quality.
In parallel efforts, some research aims to reduce translationese from human translations, and uses monolingual approaches based on text style transfer (Jalota et al., 2023), semantic parsing (Wein and Schneider, 2024) and debiasing embeddings (Dutta Chowdhury et al., 2022). Additionally, there is growing interest in leveraging human feedback to improve overall translation quality where a single metric such as COMET trained from human annotations is used as the reward model (Ramos et al., 2024; He et al., 2024). In this work we focus on improving translation quality from multiple perspectives, which is tailorable to the downstream scenario, while still being faithful to the source texts.
@@
Following neural machine translation (NMT), a new line of research started to investigate the extent to which translations (including HT and MT) contain artefacts, and how these compare to original texts and human translations.
@@
Baroni and Bernardini (2005) showed that original texts can be distinguished from human-translated texts with computational methods. Concrete textual markers, such as the frequency of function words or the use of punctuation, have been associated with this difference (Koppel and Ordan, 2011; Volansky et al., 2015). Beyond hand-crafting specific linguistic features, Pylypenko et al. (2021) find that neural architectures provide a reliable tool for distinguishing translated from original texts. They obtain state-of-the-art performance by fine-tuning multilingual BERT (Devlin et al., 2019) on the task.
@@ 3 Data
We use the parallel dataset from Toral et al. (2021), preprocessed by Ploeger et al. (2024). This dataset consists of 531 books that were originally written in English (OR) and human translated into Dutch (HT), of which 495 books for training, 5 for validation and 31 as a test set. The genres of these books vary, including literary fiction, popular fiction, non-fiction and children’s books from over 100 authors. Particularly, the test set also contains a broad range of books. [^3] A full list of author names, titles, genres and publishing years of the test set books can be found in Appendix A.1. In addition, we use monolingual data for the two baseline MT systems (see Section 5.1), consisting of a random sample of equal size to the parallel training data and disjoint from the subset used for translation detection.
@@ 4.1 Base MT Model
For HT vs OR classification, we use the monolingual Dutch data introduced in the first part of Table 1. For the other two settings, we translated a subset of the English text in the parallel data (second part of Table 1) of equal size to the monolingual training data (982,114 sentences) into Dutch using the base MT model. The resulting machine translated sentences are combined with OR texts in the monolingual data for MT vs OR classification and with HT texts in the parallel data for MT vs HT. We filter out machine translated texts that are identical to human translations. Based on the above data, we fine-tune the Dutch language model BERTje (de Vries et al., 2019) for the binary detection tasks, obtaining three different models.
Algorithm 1 Multi-perspective alignment algorithm for Naturalness and Content
@@
1:  for each step $i=0,1,\cdots,m$ do
@@ 4.3 Multi-perspective Alignment for Naturalness and Content Preservation
We introduce our method which ranks samples based on rewards that target naturalness and content preservation. This approach is inspired by recent work in text style transfer, where both meaning has to be preserved and style should be transferred (Lai et al., 2021b, a). This content vs form trade-off is similar to our situation with content preservation and naturalness. Specifically, after training a base MT model using supervised learning (Section 4.1), we further align it with human expectations in terms of naturalness and content in the form of reward learning.
@@
where $\phi$ is the parameter of the classifier. $\sigma_{t}$ is the translationese threshold, which is set to 0.5 in our experiments based on preliminary results.
@@
Where C($\cdot$) represents the COMET model and $\sigma_{t}$ represents the content threshold, which is set to 0.85 in our experiments based on preliminary results.
@@
To keep the fine-tuned model from moving too far from the base MT model, we combine the reward objective with the supervised training loss instead of using a reference model requiring large computing resources. Therefore, the final objective function of our framework consists of two components: negative log-likelihood loss in Eq. 1 and reward-based loss in Eq. 5, which are jointly formulated as
@@ 5.1 Baselines
In addition to the base MT model (Section 4.1), we include three previous methods that aim at reducing machine translationese as baselines: Tailored RR (Top-k) (Ploeger et al., 2024), automatic post-editing (APE) (Freitag et al., 2019) and Tagging (Freitag et al., 2022).
Tailored RR
is an approach that involves reranking translation candidates with a classifier that distinguishes between original and translated text. We select the Tailored RR (Top-$k$), which reranks candidates that are obtained through Top-$k$ sampling, as a baseline, since it retrieves the highest diversity in Ploeger et al. (2024).
@@ 5.2 Implementation Details
All experiments are implemented using the library HuggingFace Transformers (Wolf et al., 2020). We use the BART (Lewis et al., 2020) architecture with 6 Transformer-based (Vaswani et al., 2017) layers in both the encoder and decoder. The base MT models are trained using the AdamW optimiser (Loshchilov and Hutter, 2019) with a cosine learning rate decay, and a linear warmup of 1,000 steps. The maximum learning rate is set to 1e-4, the batch size is 256, and the gradient accumulation is 2; all reward-based models are trained with a consistent learning rate of 2e-5. We evaluate the model every 1,000 steps and use early stopping with patience 3 if the cross-entropy loss on the validation set does not decrease.
We use beam search with size 5 during inference. Since some of the training data contains instances of repeated punctuation marks, this led to the reinforcement learning method tending to optimize the model for higher rewards. Therefore, we take a simple post-processing step to remove consecutive repeated punctuation marks after the text is generated. [^4] See Appendix A.2 for post-processing examples.
@@ 5.3 Evaluation Methods
We employ three metrics to automatically calculate the content preservation of the output based on human references (and source sentences), namely BLEU (Papineni et al., 2002), COMET (Rei et al., 2020, 2022), and MetricX (Juraska et al., 2024). We use the Sacre-BLEU implementation (Post, 2018) for BLEU. Regarding the COMET family models, we use both the default model wmt22-comet-da (COMET), and the reference-free model wmt22-cometkiwi-da (KIWI) that is not used for reward learning. For MetricX, we use MetricX-24-Hybrid-XL, considering it our most important translation quality metric, since it achieved state-of-the-art performance at the WMT24 Metrics Shared Task (Freitag et al., 2024).
Figure 2: Evaluation results on the validation set under various settings. Notes: (i) The training step of 0K represents the base MT model; HM indicates the harmonic mean of classification accuracy and COMET score.
@@
We apply the translationese detection models to MT outputs and report the rate (i.e. classification accuracy) at which they are classified as the target aspect, such as OR in HT-OR, with higher rates indicating that the outputs are more human-like. Additionally, as previous studies show that translated texts are often simpler than original texts (Baker, 1993), our evaluation also covers lexical diversity. Here we report six different metrics:
@@
  • MTLD (McCarthy, 2005): evaluated sequentially as the average length of sequential word strings in a text that maintains a given TTR value. We use a threshold of 0.72, following Vanmassenhove et al. (2021). This metric has been shown to be stable across different text lengths (McCarthy and Jarvis, 2010), which is why we consider it more important a metric than TTR or Yule’s I.
  • B1 (Vanmassenhove et al., 2021): the percentage of words in the output that are in the estimated 1,000 most frequent words in a language.
@@
Figure 3: Confusion matrices for the different binary classifiers. Each row represents the results of a classifier tested on different test sets.
@@ 6.1 Initial Results
Translationese Classification
Figure 3 shows the performance confusion matrix of different binary translationese classifiers on different test sets. We can observe that for each classifier scores on the main diagonal are higher than others, with MT-OR having the highest score, followed by MT-HT and HT-OR. This is on par with the performance in similar scenarios from previous work (Pylypenko et al., 2021). While van der Werff et al. (2022) show that the distinction between the translation variants (MT and HT) is challenging, we found that human translations are even more difficult to distinguish from the original target-language texts. Interestingly, MT-OR achieves higher accuracy on the sets of HT-OR and MT-HT than their corresponding classifiers.
Machine Translation
During the alignment phase (see Section 4.3) we find that for some models the valid loss does not correlate with the naturalness aspect (i.e. classifier’s accuracy) after 1k steps: while naturalness improves, the loss on the validation set stays flat. Therefore we manually select checkpoints between the 1k and the 6k steps, and report their evaluation and loss curves in Figure 2.
The first observation is that all models achieve substantial improvement in naturalness over the first 1k steps compared to the base MT model (i.e. 0K), as reflected in the results for translationese classification (HT-OR, MT-HT and MT-OR) and lexical richness (MTLD). Although the COMET scores of some models decrease slightly, the overall score HM follows the trend of the translationese aspect. After 1k steps, MTLD scores and valid loss tend to be flat; translated language classification shows a clear improvement from 1k to 2k steps on MT-OR, a slight increase on HT-OR, and remains stable on MT-HT. For all models, although some metrics fluctuate after 2k steps, they tend to be stable overall. For the remaining experiments, we report the results of the alignment model at 5K trainingsteps.
@@
6.2 Main Results
We report the main results in Table 2, including the base MT model, the three baselines and our methods trained with both rewards: COMET for content preservation and the three different classifiers for naturalness (i.e. HT-OR, MT-HT and MT-OR).
Tailored RR achieves the highest naturalness scores (e.g. HT-OR and MTLD), but performs the worst on all translation accuracy metrics. Compared to APE, Tagging consistently performs better across the board, both in terms of content (i.e. translation accuracy) and naturalness. Additionally, we observe that using more target-original data (i.e. 4.8M vs 1M) results in lower accuracy scores but better naturalness metrics, which is consistent with our hypothesis (see Section 5.1). Overall, we observe that the three baselines underperform the base MT model in terms of translation accuracy and outperform it in most cases when it comes to naturalness metrics.
Moving to our approach, when comparing different classification rewards, the model trained with COMET & MT-HT achieves, overall, better scores than our other two models (HT-OR and MT-OR). We speculate that the rewards that foster OR do not work as well due to a mismatch between the preference of the classifier (OR) and the data in the target side of the MT training data (HT). We thus believe that such classifiers could be useful in scenarios in which the target side of the MT training data contains texts originally written in that language, which would be common in translation directions in which the target language is higher-resourced than the source language.
Overall, our best system (BM + COMET & MT-HT) achieves better naturalness scores than the base MT model (e.g. 93.3 vs 90.4 for MTLD), while even having a higher KIWI score (80.6 vs 80.4) and a lower MetricX score (2.63 vs 2.66; lower is better), two metrics that have not been used in our reward learning. Tagging attains higher naturalness scores but this comes at the price of a notable reduction in translation accuracy, as shown by KIWI (79.7 vs 80.6) and MetricX (3.05 vs 2.63).
@@ 6.2 Main Results
Table 3: Ablation study: The contribution of each reward component, where we fine-tune the base MT model using only the content reward or the naturalness reward.
@@ 6.3 Ablation Study
To assess the contribution of each reward component in our framework, we perform a set of ablation studies, the results of which are shown in Table 3. For the COMET vs COMET + classifier setting, we see higher naturalness scores in the latter in all cases for MT-HT and MT-OR (except CDU in MT-HT), as expected, while there are mixed cases in HT-OR. Also as expected, translation accuracy scores decrease when the naturalness reward is added (except COMET with MT-HT).
Compared to models using classifier-only reward, classifier + COMET results generally in better naturalness-related metrics (except PTF), but worse content-based metrics (except COMET with MT-HT). This might be due to a mismatch between the classifier’s objective and the training data (see reasoning in Section 6.2) and to complex interactions between both rewards, that would require further inspection.
@@
Table 4: Example of human-written text (source and human translation), translations of the most relevant baselines (Tagging, base MT model) and our alignment model (BM + COMET & MT-HT).